OpenStudy (anonymous):

4 ____ /160 simplify by factoring

OpenStudy (anonymous):

1/40

OpenStudy (anonymous):

is it 4 ----- ?????? 160

OpenStudy (anonymous):

\[\sqrt[4]{160}?\]

OpenStudy (anonymous):

is that better

OpenStudy (anonymous):

\[2\sqrt[4]{10}\]

OpenStudy (anonymous):

yes \[2\sqrt[4]{10}\]

OpenStudy (anonymous):

\[\sqrt[4]{160} = \sqrt[4]{2*2*2*2*2*5} = 2\sqrt[4]{10}\]